OpenStudy (anonymous):

I think its D. Solve, using the substitution method. 2x + 4y = 16 x – 7 = – 4y A. (0,-7) B. (0. 7/4) C. (9, -1/2) D. (17, -5/2)

OpenStudy (anonymous):

I'm going to add 7 to both sides of the second equation: x = 7 - 4y. Now I plug it into the first: 2 [ 7 - 4y ] + 4y = 16. 14 - 8y + 4y = 16 14 - 4y = 16 7 - 2y = 8 -1 = 2y -1/2 = y. The only choice with -1/2 is C.

OpenStudy (pinksapphire):

Yes, @TurtleMuffin is right.

OpenStudy (anonymous):

i'm always late.

OpenStudy (anonymous):

I can also multiply the entire second equation by a -1: -x + 7 = 4y. Now I recognize that my first equation already has a 4y in it! So I just plug it in directly with no distribution involved: 2x + [ -x + 7] =16 Then I simplify: x + 7 = 16 x = 9.